# STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases

### v1

@@ 3.3 Benchmark Construction
Here we detail the novel pipeline to synthesize the retrieval datasets present in Section 3.2. The core idea is to entangle relational information and textual properties, while constructing ground truth answers in an accurate and efficient way. As shown in Figure 3, the construction of the retrieval datasets generally involves four steps, and the specific processes vary depending on the characteristics of each dataset. These steps are as follows:
@@
  • Sample Relational Requirements: Initially, we sample a relation template, such as “(a product) belongs to <brand>” and instantiate it into a relational requirement, e.g., “belongs to Radio Flyer” shown in the first step of Figure 3. Each relational requirement yields a set of candidate entities that meet the requirement, i.e., products belonging to Radio Flyer. Therefore, we obtain the relational requirements and the corresponding candidates. The relation templates for the three SKBs are available in Appendix A, including 28 relation templates for the Prime SKB. Each relation template in the SKBs serves a distinct and practical purpose. For instance, the template “What is the drug that targets the genes or proteins which are expressed in <anatomy>?” is particularly valuable for precision medicine inquiries. It aids in pinpointing medications aimed at specific genetic or protein targets within defined anatomical areas, thereby contributing to the development of targeted treatments.
@@
  • Extracting Textual Properties: Subsequently, we select one of the candidate nodes that satisfies the relational requirement, which is referred to as the gold answer. We extract textual properties that align with the interests of specific roles from the textual information of the gold answer. The roles could be customers, researchers, or medical scientists, differing across the SKBs. For example, in the second step of Figure 3, we extract the phrases of functionality and user experience from a Radio Flyer product, which come from different sources in the product document. We utilize GPT-3.5-turbo-16k for the STaRK-Amazon and Claude-v2 for the other two datasets to conduct this step, where the prompts used are included in Appendix B.
@@
  • Combining Textual and Relational Information: With the textual and relational requirements, we simulate queries by fusing them using LLMs. We conduct two-stage fusion using two different LLMs, i.e., Claude-v2 and GPT-4-Turbo, to avoid bias that might arise from relying on a single LLM and ensure a more diverse set of simulated queries. Specifically, this first-stage integration is guided by various criteria, such as ensuring a natural-sounding query, adhering to the style of ArXiv searches, or aligning with specific roles. Furthermore, the second stage instructs the LLM to enrich the context and rephrase the language, thereby posing a more demanding reasoning challenge in comprehending the requirements of the query. In our example, by combining relational information “belongs to Radio Flyer” and textual information “push-along tricycle” and “fun and safe for kids”, we arrive at the final query in the third step.
@@
  • Filtering Additional Answers: In the final step, we assess whether each remaining candidate, excluding the gold answer used for extracting textual properties, meets the additional textual requirement. For the verification, we employ three Claude models. Only candidates that pass the verification across all models are included in the final ground truth answer set. Therefore, this stringent verification ensures the accuracy of the nodes in the ground truth answers. During this process, we also evaluate the accuracy of the gold nodes that pass the verification criteria. The accuracy rates STaRK-Amazon, STaRK-MAG, and STaRK-Primeare 86.6%, 98.9%, and 92.3%, respectively, demonstrating the effectiveness of our filtering approach in maintaining high-quality ground truth answers.
@@
This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs within our context. We include the complete prompts of the above steps in Appendix B.
3.4 Benchmark Data Distribution Analysis and Human Evaluation
@@
Table 4: Query diversity measurement on STaRK.
@@
  • Ratio of Relational vs. Textual Information. A key feature of our benchmark dataset is the composition of textual and relational information. Therefore, it is crucial to understand the proportionality between these two types of information. We calculate the ratio of the length of relational requirements to the length of textual requirements for each query and then average these ratios across each dataset. Intuitively, the ratio serves as an approximation of the relative distribution based on the length of requirements, and does not directly reflect the importance of each type of information in determining the final answers. As shown in Figure 5, the ratios vary across different datasets, highlighting a differing emphasis on textual versus relational information. This distribution variation presents additional challenges for retrieval systems, requiring them to adapt to the dataset characteristic and specific balance of information.
@@ 4.1 Retrieval Models
Dense Retriever. Dense Retriever finetunes a query encoder and a document encoder separately using the query answer pairs from the training dataset. We optimize the encodes using contrastive loss. Specifically, the positive pairs are constructed from the training questions and their ground truth answers, while we construct 20 hard negative answers for each query from the top false positive predictions from VSS. We use roberta-base as the base model and finetune the encodes over the entire training split for each dataset.
@@
VSS + LLM Reranker [9, 53]. This method improves the precision of the top-$v$ results from VSS by reranking them using language models, taking advantage of their advanced language understanding capabilities. For this purpose, we employ two different language models: GPT-4-turbo (gpt-4-1106-preview) and Claude-v2. We set the value of $v=20$ for the reranking process to balance between precision and computational efficiency. Specifically, we construct a prompt that instructs the language models to assign a score between 0 and 1 to a node, based on its combined textual and relational information, with certain criteria provided for rating the node. We find that this approach, which involves directly scoring nodes, performs better than asking for strict satisfaction of conditions [53].
@@ 4.2 Evaluation Metrics
Recall@$k$. Recall@$k$ measures the proportion of relevant items in the top-$k$ results. In our study, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$. This metric offers insight into the model’s ability to identify all relevant items, particularly in scenarios where missing any could be critical.
@@ 4.3 Analysis
Table 6: Main experimental results.
Dense Retriever QAGNN VSS Multi-VSS VSS+Claude2 VSS+GPT4
@@
STaRK-Amazon Hit@1 0.1529 0.2656 0.3916 0.4007 0.4266 0.4479
Hit@5 0.4793 0.5001 0.6273 0.6498 0.6746 0.7117
Recall@20 0.4449 0.5205 0.5329 0.5512 0.5376 0.5535
MRR 0.3020 0.3775 0.5035 0.5155 0.5329 0.5569
STaRK-MAG Hit@1 0.1051 0.1288 0.2908 0.2592 0.3202 0.4090
Hit@5 0.3523 0.3901 0.4961 0.5043 0.5334 0.5818
Recall@20 0.4211 0.4697 0.4836 0.5080 0.4834 0.4860
MRR 0.2134 0.2912 0.3862 0.3694 0.4129 0.4900
STaRK-Prime Hit@1 0.0446 0.0885 0.1263 0.1510 0.1611 0.1828
Hit@5 0.2185 0.2135 0.3149 0.3356 0.3582 0.3728
Recall@20 0.3013 0.2963 0.3600 0.3805 0.3598 0.3405
MRR 0.1238 0.1473 0.2141 0.2349 0.2466 0.2655
Main results. In Table 6, we report the experiment results. Firstly, we notice a significant performance gap between models trained or fine-tuned on our training datasets, such as Dense Retriever and QAGNN, and models like VSS and Multi-VSS that use embeddings generated from text-embedding-ada-002. For the dense retriever, training the encoders proves challenging when handling both textual and relational information as text, leading to training challenges. The difficulty in training a dense retriever is exacerbated by the stark differences in length between long documents and short relational pairs, making it tough for the model to learn a new knowledge base’s structure from scratch during fine-tuning. Without the advantage of pre-trained embeddings, it also struggles to align with the textual requirements. For QAGNN, it faces its own training challenges due to its high computational demands during training. Finally, the superiority of VSS and Multi-VSS demonstrates the superior pre-trained embeddings obtained from a large corpus.
@@
Significantly, the methods combining VSS with an LLM reranker consistently outperform other approaches across most metrics. Notably, reranking the top 20 predictions with an LLM leads to significant improvements in Hit@$k$ and MRR metrics compared to using VSS alone. This improvement underscores the limitations of traditional embedding methods, which lack the advanced reasoning capabilities provided by LLMs. Additionally, using GPT-4 as a reranker tends to yield better results than using Claude, further demonstrating the effectiveness of better reasoning and context understanding in improving retrieval performance.
@@
Case study. Finally, we highlight the importance of reasoning ability in achieving good retrieval results on our benchmark datasets. In Figure 6, we conduct a case study to compare the performance of VSS and VSS+Claude Reranker. Specifically, we examine a user query that requests papers from an institution on a particular topic. In this scenario, VSS fails to adequately address the relational requirement due to directly embedding the entire documents without detailed analysis. As a result, papers containing frequently repeated keywords, such as “nonlinear modeling” and “piezoelectric elements”, are mistakenly assigned high scores. However, the result improves significantly after the reranking with an LLM, where the LLM is equipped to reason the relationship between the query and the information contained in each paper, making the scores more accurately reflect the papers’ relevance to the query. This case study underscores the need for reasoning ability in grasping the complexities of queries. As embedding methods may capture some aspects of the query, they often fail to capture the detailed interaction between textual and relational information. Thus, integrating LLMs with advanced reasoning into the retrieval process can be crucial for the retrieval accuracy.
@@ 5 Conclusion
We introduce STaRK, the first benchmark designed to thoroughly evaluate the capability of retrieval systems driven by LLMs in handling semi-structured knowledge bases (SKBs). This benchmark features a diverse set of queries that are semi-structured and natural-sounding, requiring context-specific reasoning across various domains, thereby setting a new standard for assessing retrieval systems in the context of SKBs. We utilize public datasets to construct three SKBs and develop an automated, general pipeline to simulate user queries that mimic real-world scenarios. Our extensive experiments on STaRK reveal significant challenges for current models in handling both textual and relational information effectively and flexibly. Overall, STaRK offers valuable opportunities for future research to advance the field of complex and multimodal retrieval systems, where reducing retrieval latency and incorporating strong reasoning ability into the retrieval process are identified as two prospective future directions.

### v2

@@ 2.4 Benchmark Construction: Synthesized Queries
We present the novel pipeline that synthesizes the large-scale retrieval datasets. The core idea is to entangle relational information and textual properties and accurately construct ground truth answers. The construction of the retrieval datasets (Figure 3) generally involves four steps, and the specific processes vary depending on the characteristics of each dataset. These steps are as follows:
@@
  • 1) Sample Relational Requirements: Initially, we sample a relation template, such as “(a product) belongs to <brand>” and ground it into a relational requirement, e.g., “belongs to Radio Flyer” in Figure 3. Each relational requirement yields a set of candidate entities that meet the requirement, i.e., products belonging to Radio Flyer. The relation templates for the three SKBs are constructed with expert/domain knowledge, which are available in Appendix B. Each relation template in the SKBs serves a distinct and practical purpose. For instance, the template “What is the drug that targets the genes or proteins which are expressed in <anatomy>?” is particularly valuable for precision medicine inquiries. It aids in pinpointing medications aimed at specific genetic or protein targets within defined anatomical areas, thereby contributing to the development of targeted treatments.
@@
  • 2) Extracting Textual Properties: The relational requirement yields a set of satisfying candidates. Subsequently, we select one of the candidate nodes, referred to as the gold answer, and use LLMs to extract textual properties of interest from the its document. We extract the textual properties that align with the interests of specific roles, such as customers, researchers, or medical scientists, according to the specific SKB. For example, in Figure 3, we extract the phrases of functionality and user experience from a Radio Flyer product from different locations of the product document.
@@
  • 3) Combining Textual and Relational Information: After obtaining the textual and relational requirements, we synthesize them into a query using LLMs. We conduct two-stage synthesization using two different LLMs to avoid bias that might arise from relying on a single LLM and ensure a more diverse set of simulated queries. Specifically, this first-stage synthesization is guided by various criteria, such as natural sounding, adhering to the style of ArXiv searches, or aligning with specific roles. Furthermore, the second stage instructs the LLM to enrich the context and rephrase the language, demanding reasoning abilities in comprehending the requirements of the query.
@@
  • 4) Filtering Additional Answers: Finally, we employ multiple LLMs to verify if each remaining candidate, excluding the gold answer used for extracting textual properties, meets the additional textual requirement. Only candidates that pass the verification across all LLMs are included in the final ground truth answer set. This stringent verification ensures the accuracy of the nodes in the ground truth answers. To validate the precision of this filtering process, we compute the average precision of each gold answer passes the verification. The precision rates on STaRK-Amazon, STaRK-MAG, and STaRK-Prime are 86.6%, 98.9%, and 92.3%, respectively, demonstrating the effectiveness of our filtering approach in maintaining high-quality ground truth answers.
@@
This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs in our formulation. We include all of the prompts and the LLMs versions in the above steps in Appendix C.
2.5 Synthesized Data Distribution Analysis and Human Evaluation
@@
Table 4: Query diversity measurement on STaRK. See Appendix A for the metric definition.
@@
  • Ratio of Relational vs.Textual Information. A key feature of our benchmark dataset is the composition of textual and relational information. Therefore, it is crucial to understand the proportionality between these two types of information in the final queries. We calculate the ratio of relational to textual requirements for each query by comparing the word count of relational information to the word count of extracted textual properties, both of which are meta data in the constructing process. We then average these ratios across each dataset.
    Intuitively, the ratio serves as an approximation of the relative distribution based on the length of requirements, and does not directly reflect the importance of each type of information in determining the final answers. As shown in Figure 5, the ratios vary across different datasets, highlighting a differing emphasis on textual versus relational information. This distribution variation presents additional challenges for retrieval systems, requiring them to adapt to the dataset characteristic and specific balance of information.
@@
2.6 Benchmark Construction: Human-Generated Queries
To enhance the practical relevance of our benchmark dataset, we engaged 31 participants, 22 of whom are native English speakers, to generate a total of 274 queries across three SKBs. The participants were given a list of entity IDs, randomly selected from the entire entity set. They followed detailed instructions (see Appendix D) using our interactive platform to explore the SKB data. Subsequently, we manually filtered the ground truth answers, ensuring their accuracy through human verification. We then analyze the commonality and difference between synthesized and human-generated queries.
Table 6: Comparison of Human-generated and Synthesized Queries
Query Type Human-generated Query Synthesized Query
Short and Direct Red sweatshirt for proud Montreal Canadiens Suggestions for a Suunto bike mount?
Specific Author & Field Find me papers that discuss improving condenser performance authored by Stojan Hrnjak Show me papers by Seung-Hyeok Kye that discuss separability criteria.
Complex Context Help me. I am trying to diagnose a patient with persistent joint pain, and I suspect a condition where the bone is dying due to compromised blood supply, often linked to factors like steroid use, alcohol abuse, or underlying diseases - what’s the name of this sneaky bone-killing culprit? I’m experiencing joint pain accompanied by swelling, fever… I’m concerned about medications aggravating my fuzzy eyesight and potential blood clotting complications. Could you recommend treatments effective for my symptoms while minimizing these specific side effects?
Commonality. We present the statistics of the human-generated datasets in Table 3. The synthesized and human-generated queries exhibit a similar level of ambiguity, resulting in a comparable number of answers. Moreover, we find that most styles and formats of human-generated queries are represented in our synthesized query dataset. In Table 6, we demonstrate three pairs of human-generated and synthesized queries that display similarities such as short and straightforward product queries, inquiries targeting specific authors and fields, and complex contextual queries in STaRK-Prime.
Difference. We find that human-generated queries often exhibit more unique expressions compared to synthesized queries, such as "Give me a fat cross and road tire that works with my Diamondback bicycle tube" and "this sneaky bone-killing culprit" (see the last row of Table 6). This discovery suggests a future direction for our benchmark to incorporate modern and dynamic language nuances.
@@ 3.1 Retrieval Models
Dense Retriever. Dense Retriever finetunes a query encoder and a document encoder separately using the query answer pairs from the training dataset. We optimize the encoder using contrastive loss. Specifically, the positive pairs are constructed from the training questions and their ground truth answers, while we construct 20 hard negative answers for each query from the top false positive predictions from VSS. We use roberta-base as the base model and finetune the encoder over the entire training split for each dataset.
@@
VSS + LLM Reranker [9, 55]. This method improves the precision of the top-$v$ results from VSS by reranking them using language models, taking advantage of their advanced language understanding capabilities. For this purpose, we employ two different language models: GPT-4-turbo (gpt-4-1106-preview) and Claude3 (claude-3-opus). We set $v=20$ for synthesize queries and $v=10$ for human-generated queries to balance between precision and computational efficiency. Specifically, we construct a prompt that instructs the language models to assign a score between 0 and 1 to a node, based on its combined textual and relational information, with certain criteria provided for rating the node. Due to the high retrieval cost, we randomly sample 10% of the testing queries on each dataset for evaluation.
@@ 3.2 Evaluation Metrics
Recall@$k$. Recall@$k$ measures the proportion of relevant items in the top-$k$ results. For synthesized queries, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$. For human-generated queries, Recall@20 is not reported as the number of responses in the human-generated subset may exceed 20. This metric offers insight into the model’s ability to identify all relevant items, particularly in scenarios where missing any could be critical.
@@ 3.3 Analysis
Table 7: Testing results on STaRK-Syn(thesized). (*) indicates partial evaluation (10% of the testing queries) due to the high latency and cost of the methods.
Dense Retriever QAGNN VSS Multi-VSS VSS+Claude3* VSS+GPT4*
@@
STaRK-Amazon Hit@1 0.1529 0.2656 0.3916 0.4007 0.4549 0.4479
Hit@5 0.4793 0.5001 0.6273 0.6498 0.7113 0.7117
Recall@20 0.4449 0.5205 0.5329 0.5512 0.5377 0.5535
MRR 0.3020 0.3775 0.5035 0.5155 0.5591 0.5569
STaRK-MAG Hit@1 0.1051 0.1288 0.2908 0.2592 0.3654 0.4090
Hit@5 0.3523 0.3901 0.4961 0.5043 0.5317 0.5818
Recall@20 0.4211 0.4697 0.4836 0.5080 0.4836 0.4860
MRR 0.2134 0.2912 0.3862 0.3694 0.4415 0.4900
STaRK-Prime Hit@1 0.0446 0.0885 0.1263 0.1510 0.1779 0.1828
Hit@5 0.2185 0.2135 0.3149 0.3356 0.3690 0.3728
Recall@20 0.3013 0.2963 0.3600 0.3805 0.3557 0.3405
MRR 0.1238 0.1473 0.2141 0.2349 0.2627 0.2655
Table 8: Testing results on STaRK-Human(-Generated).
VSS Multi-VSS VSS+Claude3 VSS+GPT4
(ada-002) (ada-002) Reranker Reranker
STaRK-Amazon Hit@1 0.3950 0.4691 0.5602 0.5432
Hit@5 0.6419 0.7284 0.7284 0.7284
MRR 0.5265 0.5874 0.6439 0.6274
STaRK-MAG Hit@1 0.2857 0.2381 0.3810 0.3690
Hit@5 0.4167 0.4167 0.4643 0.4523
MRR 0.3581 0.3143 0.4243 0.4031
STaRK-Prime Hit@1 0.2120 0.2567 0.2844 0.2844
Hit@5 0.4037 0.4037 0.4771 0.4863
MRR 0.2984 0.3377 0.3623 0.3617
Results on Synthesized Queries. Table 7 shows the experiment results on the synthesized testing datasets. Firstly, we notice a significant performance gap between models trained or fine-tuned on our training datasets, such as Dense Retriever and QAGNN, and models like VSS and Multi-VSS that use embeddings generated from text-embedding-ada-002. For the dense retriever, training the encoders proves challenging when handling both textual and relational information as text, leading to training challenges. For QAGNN, it training challenges due to its high computational demands during training. Finally, the performance of VSS and Multi-VSS demonstrates the superior pre-trained embeddings obtained from a large corpus.
@@
Significantly, the methods combining VSS with an LLM reranker consistently outperform other approaches across most metrics. Notably, reranking the top 20 predictions with an LLM leads to significant improvements in Hit@$k$ and MRR metrics compared to using VSS alone. This improvement underscores the limitations of traditional embedding methods, which lack the advanced reasoning capabilities provided by LLMs.
@@
Results on human-generated dataset. We further evaluate the well-performing methods on human-generated queries. Table 8 presents the testing results on the human-generated dataset. Reranker methods continue to outperform other methods, confirming their superior reasoning and context understanding capabilities. Compared to synthesized datasets, the performance on human-generated queries shows highly correlated trends but generally higher metrics. This suggests that the synthesized datasets could be more difficult than the human-generated ones, offering challenging tasks and shedding light on interesting aspects of complex queries.
@@
Case study. Finally, we highlight the importance of reasoning ability in achieving good retrieval results on our benchmark datasets. In Figure 6, we conduct a case study to compare the performance of VSS and VSS+Claude3 Reranker. Specifically, we examine a user query that requests papers from an institution on a particular topic. In this scenario, VSS fails to adequately address the relational requirement due to directly embedding the entire documents without detailed analysis. As a result, papers containing frequently repeated keywords, such as “nonlinear modeling” and “piezoelectric elements”, are mistakenly assigned high scores. However, the result improves significantly after the reranking with an LLM, where the LLM is equipped to reason the relationship between the query and the information contained in each paper, making the scores more accurately reflect the papers’ relevance to the query. This case study underscores the need for reasoning ability in grasping the complexities of queries. As embedding methods may capture some aspects of the query, they often fail to capture the detailed interaction between textual and relational information. Thus, integrating LLMs with advanced reasoning into the retrieval process can be crucial for the retrieval accuracy.
4 Related Work
Unstructured QA Datasets. The line of work explores retrieving answers from a single document [33] or multiple document sources [42, 47, 22, 44, 10]. For example, SQuAD [33] focuses on extracting answers from a single document, testing model’s ability to understand and retrieve information within a specific context. For multi-document retrieval, HotpotQA [47] demands reasoning over multiple documents and includes sentence-level supporting facts, enhancing the complexity of the retrieval task. TriviaQA [22] combines questions from trivia games with evidence documents, evaluating both retrieval and reasoning over multiple sources. Moreover, some works [30, 25] leverage generated results by search engines as ground truth answers or information sources for retrieval. However, unstructured QA datasets or knowledge bases often lack the depth of relational reasoning commonly required in answering complex user queries. In contrast, STaRK contains queries demanding multi-hop relational reasoning to challenge model’s ability of handling structured information.
Structured QA Datasets. Within this category, the datasets challenge models to retrieve answers from structured knowledge bases, either knowledge graphs [3, 51, 40, 4, 5, 11, 14] or tabular data [53, 52]. For example, ComplexWebQuestions [40] challenge models’ abilities to decompose complex queries and interpret queries using entities and relationships defined in knowledge graphs. GraphQA [14] textualizes graphs by describing node attributes and edges with natural language, offering language models a way to handle KBQA. For datasets with tabular data, WikiSQL [53] and Spider [52] both address the task of converting natural language questions to SQL queries, with WikiSQL concentrating on single-table databases and Spider extending to complex multi-tables. While these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information. STaRK extends this line of work by integrating textual information into knowledge bases and queries, resulting in more comprehensive and nuanced responses with extensive insights drawn from abundant textual data.
Semi-Structured QA Datasets. Another line of work focuses on semi-structured QA with tabular and textual data. WikiTableQuestions [32] emphasizes comprehension of table structure and textual content, reflecting the semi-structured nature of many real-world datasets. TabFact [7] and HybridQA [8] combine textual and tabular data, requiring models to validate statements or answer questions using both data types. TabMCQ [20] introduces multiple-choice questions based on tables, adding complexity in understanding and reasoning over semi-structured data. However, different from this work, these existing datasets mainly focus on tables as structured sources, which may not incorporate rich relational information that naturally exists between entities. Additionally, previous attempts to combine text and tabular data through external web links or additional text columns often result in data structures that are hard to navigate. Therefore, STARK goes beyond these limitations by integrating both relational and textual data in a more cohesive manner, allowing for seamless navigation and retrieval across complex semi-structured knowledge bases. This integration supports richer relational reasoning and more comprehensive textual data handling, making STARK a more effective benchmark for evaluating the capabilities of retrieval systems in scenarios that closely mirror real-world information needs.
@@ 5 Conclusion
We introduce STaRK, the first benchmark designed to thoroughly evaluate the capability of retrieval systems driven by LLMs in handling semi-structured knowledge bases (SKBs). This benchmark features a diverse set of queries that are semi-structured and natural-sounding, requiring context-specific reasoning across various domains, thereby setting a new standard for assessing retrieval systems in the context of SKBs. We utilize public datasets to construct three SKBs and develop an automated, general pipeline to simulate user queries that mimic real-world scenarios. Moreover, we augment our datasets with human-generated queries and provide reference to the synthesized queries. Our extensive experiments on STaRK reveal significant challenges for current models in handling both textual and relational information effectively and flexibly. Overall, STaRK offers valuable opportunities for future research to advance the field of complex and multimodal retrieval systems, where reducing retrieval latency and incorporating strong reasoning ability into the retrieval process are identified as two prospective future directions.